OpenStudy (anonymous):

2. A ball is dropped from the roof of a building. The ball strikes the ground after 6 s. a) How tall, in meters, is the building? Answer: b) What is the magnitude of the ball’s velocity just before it reaches the ground? Answer:

OpenStudy (anonymous):

'free fall' question so we will use 1 of 3 kinematic equations. so where are you stuck @dwade

OpenStudy (anonymous):

pretty much which equation to use @DemolisionWolf

OpenStudy (anonymous):

awh! ok, makes sense so what I do is figure out what I know already. so like initial velocity = ? gravity = ? time = ? and once I see what I know, then I see which of the 3 equations usese the things I know. so fill in the question marks of what we 'know'

OpenStudy (anonymous):

alright so i know: initial velocity = 0 m/s gravity = 9.8 m/s time = 6 s

OpenStudy (anonymous):

easy enough. so, here are the equations to pick from v = v + at v^2 = v^2 + 2as s = s + vt + (1/2)at^2 now 's' means distance, so in our case it's height. so which of those 3 equations do we know all of the variables but s?

OpenStudy (anonymous):

v=v+at

OpenStudy (anonymous):

I probably worded my question weird asking you... but I was thinking more like this...  see how i 'check marked' the ones we have and how we are looking for s?